Line JK passes through points J(-3,11) and K(1, -3) O 7x + 2y =-1 (g) 7x+2y = 1 O 14x+4y =-1 O 14x+4y = 1
Answer #1

Line JK passes through J(-3,11) , K(1,-3)

General equation of line is y-y_{1}=m(x-x_{1})

Where 'm' is the slope

m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}

(x_{1},y_{1})=(-3,11),(x_{2},y_{2})=(1,-3)

So, slope is,

m=\frac{-3-11}{1-(-3)}=\frac{-14}{4}=\frac{-7}{2}

Substitute m=-7/2 in y-y_{1}=m(x-x_{1})

y-11=\frac{-7}{2}(x-(-3))

On multiplying with 2 on both sides

2(y-11)=-7(x+3)

2y-22=-7x-21

Shifting the terms,

7x+2y-22+21=0

7x+2y-1=0

Adding 1 on both sides

7x+2y=1

Therefore,

The answer is 7x+2y=1